# RadarVLM: A Vision-Language Model Approach for Radar Scene Understanding

### 05 Mar 2026 v3

@@ Abstract
Radar sensors provide reliable perception across adverse weather, lighting, and long-range conditions, yet existing machine learning approaches remain fragmented and task-specific, with each downstream task employing distinct architectures and training objectives. We present RadarVLM, a vision-language framework that learns unified scene-level representations through structured spatial language supervision. Leveraging the CARLA simulator with a realistic radar model, we collect over 800k radar-caption pairs across 110+ hours of simulated driving in diverse scenarios. We make two key contributions: (1) a structured caption framework encoding vehicle distributions in the radar’s native coordinate system, and (2) Spatially-Grounded CLIP (SG-CLIP) objective that replaces binary matching with continuous scene similarity, enabling fine-grained spatial reasoning. We further propose localization-aware evaluation metrics that directly assess spatial accuracy beyond traditional linguistic similarity measures. Validated on generative captioning and vehicle segmentation, SG-CLIP achieves up to 50% relative F1-score improvement over vanilla CLIP and a 21% AP gain on segmentation, demonstrating that language grounding produces spatially structured representations.
@@ 1 Introduction
Figure 1: Overview of the RadarVLM framework. Radar range-angle heatmaps are encoded by a ViT-B/16 vision encoder, while structured spatial captions are processed by a Transformer-based text encoder. Both modalities are projected into a shared embedding space where SG-CLIP aligns representations based on continuous scene similarity. The frozen encoder is subsequently validated via generative captioning (from the CLS token) and vehicle segmentation (from patch tokens).
@@
We introduce RadarVLM (Figure 1), a framework that uses language grounding to teach a radar encoder structured spatial scene understanding. To address the data challenge, we leverage the CARLA simulator integrated with a realistic radar sensor [42, 29, 2], enabling large-scale generation of diverse, well-annotated radar-caption pairs across varied autonomous driving scenarios. Our approach makes three key contributions:
@@ 2 Related Works
Radar Perception for Autonomous Driving. Existing Radar ML pipelines fall into three broad categories. Task-specific supervised methods train specialized architectures from scratch for individual perception objectives such as object detection, segmentation, or scene flow [45, 28, 50, 35, 8, 7, 16]. Cross-modal supervised methods exploit external modality supervision to generate pseudo-labels for radar training [3, 34, 8], while SSL approaches use radar-to-radar and radar-to-vision to pretrain radar embeddings [14]. Multi-sensor fusion architectures combine radar with complementary sensors [7, 4]. Across all these paradigms, learned representations remain tied to specific tasks and architectures, and none provide the spatial semantics needed for relational scene understanding.
Contrastive Loss Variants. Several works have extended CLIP’s binary matching with softer supervision, including soft-label relaxations [13], supervised contrastive clustering [19], prototype-based objectives [26], ranking-consistent losses [51], and sigmoid-based or multi-task variants [49, 31]. Closer to our setting, mmCLIP [5] demonstrates the viability of CLIP-style alignment for non-visual sensing by grounding mmWave signals in language. However, none of these address the continuous spatial similarity structure between radar scenes inherent to autonomous driving. Our SG-CLIP objective fills this gap through soft similarity targets derived from per-cell vehicle count overlap.
Radar-Language Alignment. The most closely related works to ours are early explorations of radar-language grounding, this work [37] applies CLIP-style contrastive pretraining between radar spectrograms and text, showing preliminary evidence of improved downstream scene parsing. RadarLLM [22] connects millimeter-wave point cloud sequences to LLMs for human motion understanding. While these works establish the feasibility of radar-language alignment, they rely on generic scene-level descriptions that lack explicit spatial grounding. In contrast, we address both limitations directly, and enables fine-grained spatial learning rather than coarse scene-level matching.
@@ 3 Dataset Curation
Contrastive VLM pre-training requires large-scale paired data to learn transferable cross-modal representations [38]; however, the scale of real-world annotated radar datasets is not sufficient, motivating our use of simulation. We leverage the CARLA simulator [11], a well-established platform for autonomous driving research, which provides perfect access to ground truth data for all actors in the scene. This simulator-based approach enables us to collect radar data with precise spatial annotations at scale. Since CARLA’s radar sensor is highly simplified, we use an open-source implementation [42, 29, 2] which models the propagation behavior of automotive radars very accurately.
@@
Data collection and statistics: We collect diverse driving scenarios across CARLA’s urban, highway, and intersection environments under varying traffic densities. For each radar frame, we extract the complete ground truth state of all actors and filter based on the ego vehicle’s location, retaining only objects within the radar’s sensing range. The traffic scenarios range from sparse (1 to 2 vehicles visible) to dense (10+ vehicles), ensuring the model learns across varying complexity levels. Figure 4(a) shows the distribution of the number of vehicles per scene across the entire dataset, while Figure 4(b) illustrates the lane-wise spatial distribution of collected vehicles. This dataset is, to our knowledge, the first large-scale radar dataset with structured, spatially-grounded natural language descriptions, and we will release it to facilitate future research.
@@
4 RadarVLM: Architecture and Approach
Our framework, RadarVLM, comprises two key components: a CLIP-based [38, 10] vision encoder for radar heatmaps and a GPT-2-like Transformer [39, 43] text encoder for captions, unified by the Spatially-Grounded Contrastive Learning (SG-CLIP) objective that accounts for continuous scene similarity. While contrastive pretraining aligns radar and text representations at the scene level, it alone cannot confirm that the learned representations are genuinely spatially grounded.
To rigorously validate spatial grounding, we extend RadarVLM on two complementary downstream tasks of generative captioning and segmentation (see Figure 1). Generative captioning interrogates whether the global CLS token encodes sufficient structured semantic information to decode precise vehicle distributions by distance bin and angular sector into accurate natural-language descriptions. Vehicle segmentation directly interrogates the spatial organization of the encoder’s patch-level features, probing whether spatial structure is preserved in the encoder’s internal representations at the pixel level without any linguistic mediation. In both tasks, the vision encoder is kept frozen. A model that succeeds at both has demonstrated spatial grounding at two distinct levels of abstraction, providing strong evidence that SG-CLIP produces representations that are spatially grounded throughout.
@@ 4.2 Spatially-Grounded Contrastive Learning
Soft target matrix: We construct a soft target matrix $\mathbf{T}^{\text{soft}}$ for each batch (size $N$) by row-normalizing the pairwise similarities. The normalization ensures a valid probability distribution.
@@ 5 Experiments
Training Setup: All three training stages operate on the same 80/20 train-test split and use the AdamW optimizer, training until validation loss converges. For pretraining, we use a batch size of 160, a cosine learning-rate schedule, and an initial learning rate of $10^{-5}$. The ClipCap captioning decoder is trained with a batch size of 32 and a learning rate of $10^{-4}$, while the segmentation decoder uses a batch size of 512 and a learning rate of $10^{-3}$. All experiments are conducted on a cluster of four NVIDIA A100 GPUs. Wall-clock training time is approximately one day for all tasks.
SG-CLIP produces radar representations that improve both structured scene description and pixel-level spatial localization over standard CLIP-style pre-training. We first qualitatively verify that SG-CLIP concentrates encoder attention on vehicle-occupied regions (Section 5.1), and then quantitatively demonstrate this through: (1) generative captioning assessed with localization-aware metrics, where SG-CLIP achieves up to 50% relative F1-score improvement at long range over vanilla CLIP [38] (Section 5.3), and (2) vehicle segmentation, where SG-CLIP features yield a 5% IoU gain and 21% AP gain over vanilla CLIP [38] and UNet [41, 52] based segmentations (Section 5.4).
@@ 5.3 Generative Captioning Performance
Second, softer similarity kernels (lower $\alpha$) yield stronger captioning performance. Note that as $\alpha$ increases, the Gaussian kernel concentrates all mass on exact matches, recovering vanilla CLIP’s binary labels. The $\alpha=1.0$ variant achieves the best F1 at three of four distance bins, confirming that distributing the gradient signal across partially similar scenes improves fine-grained spatial reasoning over near-binary matching. Third, while performance degrades with distance for all methods, as expected given weaker vehicle reflections at range, the widening performance gap between SG-CLIP and vanilla CLIP indicates that graded supervision specifically addresses this long-range challenge. Figure 6 visualizes the per-cell F1 distribution, showing that SG-CLIP variants achieve more spatially uniform accuracy across lane sectors compared to vanilla CLIP.
@@ 5.4 Vehicle Segmentation
Several trends stand out. First, all language-pretrained encoders substantially outperform the pre-trained U-Net [41], which is trained on the dataset (0.489 IoU), despite using only a lightweight decoder on frozen features. This confirms that contrastive pre-training with VLMs transfers meaningful spatial structure to patch-level representations, even though it operates only on the global CLS token. Second, the optimal bandwidth ($\alpha=4.0$) differs from the captioning-optimal $\alpha=1.0$, consistent with the different abstraction levels each task probes: segmentation benefits from moderately hard contrastive objectives that sharpen spatial boundaries, while captioning benefits from softer kernels that encourage fine-grained distributional reasoning. Third, the 21% AP improvement from vanilla CLIP to SG-CLIP indicates that graded spatial supervision substantially improves the encoder’s ability to discriminate vehicle-occupied regions across confidence thresholds. Figure 9 confirms this trend, showing SG-CLIP variants dominating across all thresholds.

### 12 Mar 2026 v4

# RLM: A Vision-Language Model Approach for Radar Scene Understanding

@@ Abstract
Radar sensors enable reliable perception across adverse weather, low-visibility, and long-range conditions, making them critical for robust autonomous driving systems. However, current machine learning approaches for radar remain fragmented, with downstream tasks relying on separate, task‑specific architectures and training objectives. We present Radar Language Model (RLM), a vision-language framework that learns unified scene-level representations through structured spatial language supervision. Leveraging the CARLA simulator with a realistic radar model, we collect over 800k radar-caption pairs across 110+ hours of simulated driving in diverse scenarios. We make two key contributions: (1) a structured caption framework encoding vehicle distributions in the radar’s native coordinate system, and (2) Spatially-Grounded CLIP (SG-CLIP) objective that replaces binary matching with continuous scene similarity, enabling fine-grained spatial reasoning. We further propose localization-aware evaluation metrics. Validated on generative captioning and vehicle segmentation, SG-CLIP consistently outperforms Vanilla CLIP, achieving upto 50% relative F1-score improvement and a 21% AP gain on segmentation, demonstrating that language grounding produces spatially structured representations.
@@ 1 Introduction
Figure 1: Overview of the Radar Language Model (RLM) framework. Radar range-angle heatmaps are encoded by a ViT-B/16 vision encoder, while structured spatial captions are processed by a Transformer-based text encoder. Both modalities are projected into a shared embedding space where SG-CLIP aligns representations based on continuous scene similarity. The frozen encoder is subsequently validated via generative captioning (from the CLS token) and vehicle segmentation (from patch tokens).
@@
We introduce Radar Language Model (RLM) (Figure 1), a framework that uses language grounding to teach a radar encoder structured spatial scene understanding. To address the data challenge, we leverage the CARLA simulator integrated with a realistic radar sensor [44, 30, 2], enabling large-scale generation of diverse, well-annotated radar-caption pairs across varied autonomous driving scenarios. Our approach makes three key contributions:
@@ 2 Related Works
Radar Perception for Autonomous Driving. Existing Radar ML pipelines fall into three broad categories. Task-specific supervised methods train specialized architectures from scratch for individual perception objectives such as object detection, segmentation, or scene flow [47, 29, 52, 36, 9, 8, 17]. Cross-modal supervised methods exploit external modality supervision to generate pseudo-labels for radar training [3, 35, 9], while SSL approaches use radar-to-radar and radar-to-vision to pretrain radar embeddings [15]. Multi-sensor fusion architectures combine radar with complementary sensors [8, 4]. Across all these paradigms, learned representations remain tied to specific tasks and architectures, and none of these provide the spatial semantics needed for relational scene understanding.
Contrastive Loss Variants. Several works have extended CLIP’s binary matching with softer supervision, including soft-label relaxations [14], supervised contrastive clustering [20], prototype-based objectives [27], ranking-consistent losses [53], and sigmoid-based or multi-task variants [51, 32]. Closer to our setting, mmCLIP [5] demonstrates the viability of CLIP-style alignment for non-visual sensing by grounding mmWave signals in language. Our work shares the broad motivation of replacing binary targets, but differs in how similarity is defined: prior methods derive soft targets from generic embedding distances, label hierarchies, whereas SG-CLIP constructs similarity directly from per-cell vehicle count overlap in an ego-centric spatial grid, unique to radar driving scenes.
Radar-Language Alignment. The most closely related works to ours are early explorations of radar-language grounding, this work [38] applies CLIP-style contrastive pretraining between radar spectrograms and text, showing preliminary evidence of improved downstream scene parsing. RadarLLM [23] connects millimeter-wave point cloud sequences to LLMs for human motion understanding. While these works establish the feasibility of radar-language alignment, they rely on generic scene-level descriptions that lack explicit spatial grounding. In contrast, we address both limitations directly and enable fine-grained spatial learning rather than coarse scene-level matching.
@@ 3 Dataset Curation
Contrastive VLM pre-training requires large-scale paired data to learn transferable cross-modal representations [39]; however, the scale of real-world annotated radar datasets is not sufficient, motivating our use of simulation. We leverage the CARLA simulator [12], a well-established platform for autonomous driving research, which provides perfect access to ground truth data for all actors in the scene. This simulator-based approach enables us to collect radar data with precise spatial annotations at scale. Since CARLA’s radar sensor is highly simplified, we use an open-source implementation [44, 30, 2] which models the propagation behavior of automotive radars very accurately. We present results on this simulated dataset here; to assess generalization beyond simulation, we additionally experiment with the real-world datasets [4, 43] in the supplementary material.
@@
Data collection and statistics: We collected diverse driving scenarios across CARLA’s urban, highway, and intersection environments under varying traffic densities. For each radar frame, we extract the complete ground truth state of all actors and filter based on the ego vehicle’s location, retaining only objects within the radar’s sensing range. The traffic scenarios range from sparse to dense vehicles, ensuring the model learns across varying complexity levels. Figure 4(a) shows the distribution of the number of vehicles per scene across the entire dataset, while Figure 4(b) illustrates the lane-wise spatial distribution of collected vehicles. This dataset is, to our knowledge, the first large-scale radar dataset with structured, spatially-grounded natural language descriptions, and we will release it to facilitate future research.
@@
4 RLM: Architecture and Approach
Our framework, RLM, comprises two key components: a CLIP-based [39, 11] vision encoder for radar heatmaps and a GPT-2-like Transformer [40, 45] text encoder for captions, unified by the Spatially-Grounded Contrastive Learning (SG-CLIP) objective that accounts for continuous scene similarity. While contrastive pretraining aligns radar and text representations at the scene level, it alone cannot confirm that the learned representations are genuinely spatially grounded.
To rigorously validate spatial grounding, we extend RLM on two complementary downstream tasks of generative captioning and segmentation (see Figure 1). Generative captioning interrogates whether the global CLS token encodes sufficient structured semantic information to decode precise vehicle distributions by distance bin and angular sector into accurate natural-language descriptions. Vehicle segmentation directly interrogates the spatial organization of the encoder’s patch-level features, probing whether spatial structure is preserved in the encoder’s internal representations at the pixel level without any linguistic mediation. In both tasks, the vision encoder is kept frozen. A model that succeeds at both has demonstrated spatial grounding at two distinct levels of abstraction, providing strong evidence that SG-CLIP produces representations that are spatially grounded throughout.
@@ 4.2 Spatially-Grounded Contrastive Learning
Soft target matrix: We construct a soft target matrix $\mathbf{T}^{\text{soft}}$ for each batch (size $N$) by row-normalizing the pairwise similarities. The normalization ensures a valid probability distribution and prevents gradient explosion during training.
@@ 5 Experiments
Training Setup: All three training stages operate on the same 80/20 train-test split, use the AdamW optimizer, and training proceeds until validation loss converges. For pretraining, we use a batch size of 160, a cosine learning-rate schedule, and an initial learning rate of $10^{-5}$. The ClipCap captioning decoder is trained with a batch size of 32 and a learning rate of $10^{-4}$, while the segmentation decoder uses a batch size of 512 and a learning rate of $10^{-3}$. All experiments are conducted on a cluster of four NVIDIA A100 GPUs. Wall-clock training time is approximately one day for all tasks.
SG-CLIP produces radar representations that improve both structured scene description and pixel-level spatial localization over standard CLIP-style pre-training. We first qualitatively verify that SG-CLIP concentrates encoder attention on vehicle-occupied regions (Section 5.1), and then quantitatively demonstrate this through: (1) generative captioning assessed with localization-aware metrics, where SG-CLIP consistently improves at long range over vanilla CLIP [39] (Section 5.3), and (2) vehicle segmentation, where SG-CLIP features yield a 5% IoU gain and 21% AP gain over vanilla CLIP [39] and UNet [42, 54] based segmentations (Section 5.4).
@@ 5.3 Generative Captioning Performance
Second, softer similarity kernels (lower $\alpha$) yield stronger captioning performance. Note that as $\alpha$ increases, the Gaussian kernel concentrates all mass on exact matches, recovering vanilla CLIP’s binary labels. The $\alpha=1.0$ variant achieves the best F1 at three of four distance bins, confirming that distributing the gradient signal across partially similar scenes improves fine-grained spatial reasoning over near-binary matching.
Third, while performance degrades with distance for all methods, as expected given weaker vehicle reflections at range, the widening performance gap between SG-CLIP and vanilla CLIP indicates that graded supervision specifically addresses this long-range challenge. Figure 6 visualizes the per-cell F1 distribution, showing that SG-CLIP variants achieve more spatially uniform accuracy across lane sectors compared to vanilla CLIP.
@@ 5.4 Vehicle Segmentation
Several trends stand out. First, all language-pretrained encoders substantially outperform the pre-trained U-Net [42], which is trained on the dataset (0.489 IoU), despite using only a lightweight decoder on frozen features. This confirms that contrastive pre-training with VLMs transfers meaningful spatial structure to patch-level representations, even though it operates only on the global CLS token. Second, the optimal bandwidth ($\alpha=4.0$) differs from the captioning-optimal ($\alpha=1.0$), consistent with the different abstraction levels each task probes: segmentation benefits from moderately hard contrastive objectives that sharpen spatial boundaries, while captioning benefits from softer kernels that encourage fine-grained distributional reasoning. Third, the 21% AP improvement from vanilla CLIP to SG-CLIP indicates that graded spatial supervision substantially improves the encoder’s ability to discriminate vehicle-occupied regions across confidence thresholds. Figure 9 confirms this trend, showing SG-CLIP variants dominating across all thresholds.